If consumption is C=100+0.75Yd

Taxes is T=50+0.5Y

Export is X=200

Import is M=50+0.25Y

Government spending is G=150

Investment is I=200. Y is domestic income and Y is private disposable income. Determine the surplus or deficit in the government budget at equilibrium


1
Expert's answer
2021-05-16T17:49:28-0400

C=100+0.75Yd

T=50+0.5Y

X=200

M=50+0.25Y

G=150

I=200.

Now, 

"C = 100+0.75Yd"

"= 100+0.75(Y-T)"

  "= 100+0.75[Y-(50+0.5Y)]"

"=100+0.75(Y-50-0.5Y)"

"= 100+0.75(0.5Y-50)"


  "Y = C+I+G+ (X-M)"

"=> Y = 100+0.75(0.5Y-50) + 200+ 150 + ( 200 - 50 - 0.25Y)"

"=> Y = 100 + 0.375Y - 37.5 + 200+ 150 + 150 - 0.25Y"

"=> Y - 0.375Y + 0.25Y = 562.5"

"=> 0.875Y = 562.5"

"=> Y = 642.86"

Hence "Y = 642.86"

Therefore, "T = 50+0.5Y"

        "T = 50+(0.5\times642.86)"

         "T = 371.43"

Budget"= T - G"

     "= 371.43 - 150"

      "= 221.43"

Hence there is the surplus in the government budget at equilibrium as Government spending is positive.